A TRIANGLE HAS AN AREA OF 128 SQUARE INCHES. THE BASE OF THE TRIANGLE MEASURE 16 INCHES. FIND THE HEIGHT OF THE TRIANGLE IN INCHES.
Answer by checkley77(12844)

A=BH/2
128=16H/2
16H=2*128
16H=256
H=256/16
H=16 IN. ANS. FOR THE HEIGHT.